OpenStudy (ksaimouli):

Let f be the continuous function defined on [−4, 3] whose graph, consisting of three line segments and a semicircle centered at the origin, is given above. Let g

OpenStudy (ksaimouli):

\[g(x)=\int\limits_{1}^{x}f(t)dt\]

OpenStudy (ksaimouli):

find g(-2)

OpenStudy (ksaimouli):



OpenStudy (ksaimouli):

@hartnn

OpenStudy (anonymous):

Solve this I have a container with 4,200 milliliters of water. How many litres are in my container?

OpenStudy (ksaimouli):

i need to find g(-2)

OpenStudy (anonymous):

the equation is the area of a triangle minus the area of a semi circle... do you see that?

OpenStudy (anonymous):

the traingle has height = 3 and base = 1, so A(triangle)=1/2 baseXheight = (1/2)(1)(3) = 3/2

OpenStudy (anonymous):

the semi circle has area of (1/2)PiR^2 where R=1, That means the area = Pi/2

OpenStudy (ksaimouli):

g(2)

OpenStudy (anonymous):

\[\int\limits_{1}^{-2} = -\int\limits_{-2}^{1} \]

OpenStudy (ksaimouli):

yup

OpenStudy (ksaimouli):

i got it thx what about g(2)

OpenStudy (ksaimouli):

usind distance formula?

OpenStudy (anonymous):

Im thinking there might be an easier way,

OpenStudy (ksaimouli):

ok

OpenStudy (anonymous):

\[(\frac{ -1-1 }{ 3-1 })(1)\] will give you the height of the triangle

OpenStudy (ksaimouli):

how did u get that

OpenStudy (anonymous):

then the base is 1

OpenStudy (ksaimouli):

@Edutopia how did u get that ?

OpenStudy (ksaimouli):

@Edutopia

OpenStudy (anonymous):

yeah, it should be (-1-0)/(3-1), my mistake: that is the slope which is the change in y with respect to x, so as x change one (the distance between 1 and 2) y will change -1/2

OpenStudy (ksaimouli):

what points did u choose

OpenStudy (anonymous):

A(triangle)=(-1/2)(1)(1/2), i used the two given points you have above in your graph

OpenStudy (ksaimouli):

ok

OpenStudy (ksaimouli):

is that same for g'(-3)